OpenStudy (anonymous):

The legs of a right triangle have lengths of 4 and 5 units. Find the length of the hypotenuse

OpenStudy (rmrjr22):

a^2 + b^2 = c^2

OpenStudy (rmrjr22):

c^2 = hypotnuse

OpenStudy (jhannybean):

 \(\large c^2 = (4)^2 + (5)^2\) So \(\large c = \sqrt{4^2 +5^2 }\)

OpenStudy (rmrjr22):

yup

OpenStudy (rmrjr22):

\[\sqrt{41}\]

OpenStudy (jhannybean):

\[\large \color{red}\checkmark\]